OpenStudy (anonymous):

xy^4 ----- how do you do this? 6x^2y^4

OpenStudy (anonymous):

\[\frac{y^4}{y^4}=1\]

OpenStudy (anonymous):

\[\frac{x}{x^2}=\frac{1}{x}\]

OpenStudy (anonymous):

leaving \[\frac{1}{6x}\]

OpenStudy (anonymous):

i dont get it